OpenStudy (anonymous):

sqrt20 + sqrt40. solve

OpenStudy (anonymous):

2sqrt5 + 2sqrt10 = 2(sqrt5+sqrt10)

OpenStudy (anonymous):

\[\sqrt{20}+\sqrt{40}\]\[\sqrt{4\times5}+\sqrt{4\times10}\]\[2\sqrt{5}+2\sqrt{10}\]\[2(\sqrt{5}+\sqrt{10})\]